find v if T=427 degrees celcius and P=39MPa and no qaulity is given fluid is water. clear explaination and work should be shown
Answer #1

\textbf{Given data}\rightarrow

T=427\: \: ^{o}\textup{C}

P=39\textup{ MPa}

\textbf{Solution}\rightarrow

\textup{From superheated steam table properties 35 MPa and 427 }^{o}\textup{C}

T\textup{ in }^{o}\textup{C} v\textup{ in m}^{3}\textup{/kg}
420 0.00308380
427 v
430 0.00378000

\textup{By interpolation}

\frac{430-420}{0.0037800-0.00308380}=\frac{430-427}{0.0037800-v_{1}}

v_{1}=0.00357114\textup{ m}^{3}\textup{/kg}

\textup{From superheated steam table properties 40 MPa and 427 }^{o}\textup{C}

T\textup{ in }^{o}\textup{C} v\textup{ in m}^{3}\textup{/kg}
420 0.00236010
427 v
430 0.00274370

\textup{By interpolation}

\frac{430-420}{0.00274370-0.00236010}=\frac{430-427}{0.00274370-v_{2}}

v_{2}=0.00262862\textup{ m}^{3}\textup{/kg}

\textup{From superheated steam table properties 39 MPa and 427 }^{o}\textup{C}

P\textup{ in MPa} v\textup{ in m}^{3}\textup{/kg}
35 v_{1}=0.00357114\textup{ m}^{3}\textup{/kg}
39 v
40 v_{2}=0.00262862\textup{ m}^{3}\textup{/kg}

\textup{By interpolation}

\frac{40-35}{0.00262862-0.00357114}=\frac{40-39}{0.00262862-v}

v=0.00281712\textup{ m}^{3}\textup{/kg}